# Your Finetuned Large Language Model is Already a Powerful Out-of-distribution Detector

### v1

@@ 5.2 Near OOD Detection
Yes Podolskiy et al. (2021) 0.998 0.994 0.008
SNIPS No Gangal et al. (2020) 0.955 0.903 0.192
Jin et al. (2022) 0.963 0.910 0.145
@@
Yes Podolskiy et al. (2021) 0.978 0.933 0.120
CLINC150 No Gangal et al. (2020) 0.883 0.677 0.463
Jin et al. (2022) 0.902 0.703 0.417
@@
Yes Podolskiy et al. (2021) 0.982 0.939 0.092
Table 2: Results for near OOD detection. Since the experimental configurations in the studies by Gangal et al. (2020), Podolskiy et al. (2021), and Jin et al. (2022) differ, we have replicated their methods and aligned the dataset splitting for consistency.
5.3 Spam Detection
Given that the concept of unsupervised OOD detection aligns closely with spam detection, we evaluate our method using the spam detection benchmark introduced by Labonne & Moran (2023). This benchmark includes four specific spam detection datasets: Ling-Spam Dataset (Sakkis et al., 2003), SMS Spam Collection (Almeida et al., 2011), SpamAssassin Public Corpus, and Enron Email Dataset (Metsis et al., 2006). It compares the performance of both traditional and deep learning-based binary classifiers. Our method, being rooted in OOD detection, requires only the non-spam (ham) data for finetuning the LLM. Table 3 indicates that our method, without any spam data, can still reliably identify spam. Furthermore, when spam data is available, we can finetune the OOD proxy using this data and apply the likelihood ratio between the two finetuned LLMs. This approach achieves performance that is competitive with the SOTA in spam detection.
5.4 OOD Question Detection in QA Systems
We test the effectiveness of the criterion we introduced in Section 4 within a QA scenario. Here, we employ MetaMath (Yu et al., 2023), which is a Llama2 model finetuned on mathematics data for solving mathematical problems. The objective in this QA context is to identify OOD questions that originate from domains outside the model’s expertise. Filtering out OOD questions enhances the system’s robustness. For this evaluation, we designate the test sets of the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) datasets as In-D and use SQUAD (Rajpurkar et al., 2018), BoolQ (Clark et al., 2019), and PIQA (Bisk et al., 2020) as OOD datasets. Table 4 indicates that the criterion $S_{a|q}$ consistently outperforms the other evaluated criteria, with all its AUROC values exceeding 0.5. This demonstrates its effectiveness in detecting OOD. Notably, $S_{q}$ exhibits subpar performance in most scenarios, underscoring the necessity of our proposed approach that generates an answer and formulates the criterion based on the question-and-answer pair.
@@ 5.4 OOD Question Detection in QA Systems
Table 3: Results of spam detection. This table presents a condensed version; for the complete details, please see Table 6 in Appendix.
@@
GSM8K SQUAD 2.0 7B $S_{q}$ 0.1116 0.0546 0.9894
$S_{a}$ 0.5463 0.0979 0.9947
$S_{q,a}$ 0.5363 0.0959 0.9924
$S_{a|q}$ 0.6877 0.1376 0.9704
BoolQ 7B $S_{q}$ 0.0538 0.1618 0.9955
$S_{a}$ 0.5045 0.2616 0.9932
$S_{q,a}$ 0.4797 0.2536 0.9879
$S_{a|q}$ 0.7156 0.4041 0.9310
PIQA 7B $S_{q}$ 0.9762 0.9779 0.0735
$S_{a}$ 0.9612 0.9746 0.0569
$S_{q,a}$ 0.9975 0.9983 0.0038
$S_{a|q}$ 0.9944 0.9944 0.0099
MATH SQUAD 2.0 7B $S_{q}$ 0.2139 0.2304 0.9474
@@
Table 4: Outcomes of OOD question detection in QA settings. This table presents a condensed version; for the complete details, please see Table 7 in Appendix.
@@ 6 Discussions
Our far OOD detection experiments in Table 1 show that the AUROC for likelihood-based OOD detection methods can be exceptionally low. This finding echoes the phenomenon highlighted by Nalisnick et al. (2018), where language OOD data may unexpectedly exhibit higher likelihood values. Upon examining the characteristics of the datasets involved, it becomes apparent that the texts from the 20 Newsgroups (20NG) (Lang, 1995) dataset are significantly longer than those from the comparative OOD datasets, such as SST-2 and RTE, especially in instances where the AUROC is notably low. This observation reveals a tendency among language models to assign higher likelihoods to shorter sentences, irrespective of their actual semantic content. It highlights the crucial importance of adopting the likelihood ratio rather than solely relying on raw likelihood values to enhance the accuracy of OOD detection.
@@
In the near OOD detection experiments presented in Table 2, Podolskiy et al. (2021)’s approach outperforms competing methods significantly. This superior performance may be attributed to the distinct class distribution across datasets. Specifically, the CLINC150 dataset comprises 150 In-D classes, in stark contrast to the SNIPS and ROSTD datasets, which offer a mere 7 (with only 5 for In-D) and 13 In-D classes, respectively. The substantially greater number of In-D classes in CLINC150 compared to the other datasets likely enhances Podolskiy et al. (2021)’s method’s ability to leverage the extensive class label information, thus yielding improved OOD detection results.
7 Related Work
Building upon the OOD detection method using likelihood ratios introduced by Ren et al. (2019), Gangal et al. (2020) suggested employing the likelihood ratio between two LSTM language models. In their approach, one model functions as a “background model” representing OOD data and is trained on random combinations from the vocabulary.
Jin et al. (2022) used the likelihood ratio between a pretrained GPT-2 and a finetuned version of GPT-2 as a baseline to compare with their proposed contrastive learning-based OOD detection method. They showed that their proposed method can out perform the likelihood ratio based method. However, they only used GPT-2 in their likelihood ratio baseline, which is very small both in terms of training data and the model size by today standard.
In comparison, our study provides a more comprehensive analysis of likelihood ratio between base models and fine-tuned models using much larger LLMs. We show that leveraging these more advanced LLMs, likelihood ratio significantly outperforms results from Jin et al. (2022) in OOD detection. Additionally, in the current era, accessing and sharing both pretrained and finetuned LLMs has become considerably easier. Using likelihood ratio in LLMs for OOD detection is easy, accessible, and very effective.
@@ 8 Concluding Remarks
Appendix A Discussion (Continued)
A.1 The cases that LH is better than LR
In the spam detection experiments detailed in Table 3, particularly with data from the SMS and SpamAssassin datasets, we observe that without spam data, the likelihood (LH) method outperforms the likelihood ratio (LR). While LR generally shows superior and more consistent performance across most experiments - as LH can exhibit extremely poor performance in certain cases, a point elaborated in Section 6 - there are specific instances where LH is more effective.
A.2 A Qualitative Analysis on QA OOD Detection
Figure 3 demonstrates sample QA pairs generated by MetaMath, accompanied by their associated criteria: $S_{q}$, $S_{a}$, $S_{q,a}$, and $S_{a|q}$, which are introduced in Section 4. A higher score for a given criterion suggests that the QA pair is more likely to be OOD. Example (a) is a classic In-D instance, presenting a mathematical problem paired with a logical solution that culminates in a quantitative answer. Example (b) poses a conventional question paired with an irrelevant quantitative answer under the guise of mathematical logic. Example (c) is an undesirable case where MetaMath repeats sentences without providing a substantial answer. Examples (d) and (f) display standard questions with appropriate responses from MetaMath. In example (e), MetaMath attempts to address a general question through mathematical reasoning but does not succeed.
The criterion $S_{q}$ for example (a) exhibits the highest value among all $S_{q}$ scores, indicating that $S_{q}$ may not effectively detect OOD questions. This observation is consistent with our analysis in Section 4 and our discussion in Section 6. As for the most effective score, $S_{a|q}$, we observe that when MetaMath provides an appropriate response to a general question, $S_{a|q}$ is able to effectively flag it as an OOD question, as demonstrated in examples (d) and (f). This supports the hypothesis that a typical answer generated from a conventional question is more likely to receive a higher likelihood in Llama - a model oriented towards natural language - compared to MetaMath, which is specialized in mathematical problems.
[…]
Figure 3: Example question-answer sets produced by MetaMath-7B, with the applicable criteria displayed above each pair. The criteria are represented on a logarithmic scale.
Appendix B Finetuning Details
For finetuning, with the exception of MetaMath (Yu et al., 2023), which is comprehensively finetuned from the Llama-2 model, we employ LoRA (Low-Rank Adaptation) as described by Hu et al. (2021) for adjusting the models. Our aim is to demonstrate the general applicability of our method; therefore, all our LoRA finetuning follows a uniform parameter configuration. In accordance with prevalent practices, we finetune the parameters in the Q (query) and V (value) projections. We employ a LoRA rank of 16, a LoRA alpha of 32, and a LoRA dropout rate of 0.05. Given the varying sizes of datasets, we adjust the learning rate between $10^{-3}$ and $10^{-4}$, and the training typically spans 10 to 20 epochs. We select the model checkpoint with the best evaluation loss as our finetuned model. For detailed information, please refer to the code provided.
Appendix C Full Experiment Results
Full experiment results are detailed below.
In-D OOD Method AUROC $\uparrow$ AUPR (OOD) $\uparrow$ FPR95 $\downarrow$
20NG SST-2 Zhou et al. (2021) 0.978 0.865 0.015
CE (Hendrycks & Gimpel, 2016) 0.981 0.942 0.087
TAPT (Gururangan et al., 2020) 0.981 0.939 0.088
SupCon (Khosla et al., 2020) 0.980 0.943 0.094
Uppaal et al. (2023) 1.000 0.999 0.000
Llama-7B LH 0.008 0.541 0.999
Llama-7B LR 1.000 1.000 0.000
Mistral-7B LH 0.008 0.541 1.000
Mistral-7B LR 0.995 0.999 0.009
Llama-13B LH 0.009 0.541 1.000
Llama-13B LR 1.000 1.000 0.000
MNLI Zhou et al. (2021) 0.964 0.978 0.224
CE (Hendrycks & Gimpel, 2016) 0.968 0.989 0.166
TAPT (Gururangan et al., 2020) 0.964 0.988 0.175
SupCon (Khosla et al., 2020) 0.970 0.990 0.156
Uppaal et al. (2023) 1.000 1.000 0.000
Llama-7B LH 0.020 0.119 0.999
Llama-7B LR 1.000 1.000 0.001
Mistral-7B LH 0.024 0.119 0.999
Mistral-7B LR 0.996 0.996 0.008
Llama-13B LH 0.024 0.119 0.998
Llama-13B LR 1.000 1.000 0.000
RTE Zhou et al. (2021) 0.956 0.860 0.312
CE (Hendrycks & Gimpel, 2016) 0.945 0.902 0.285
TAPT (Gururangan et al., 2020) 0.919 0.869 0.352
SupCon (Khosla et al., 2020) 0.952 0.914 0.248
Uppaal et al. (2023) 1.000 0.999 0.000
Llama-7B LH 0.063 0.443 0.998
Llama-7B LR 1.000 1.000 0.001
Mistral-7B LH 0.074 0.446 0.998
Mistral-7B LR 0.997 0.999 0.006
Llama-13B LH 0.070 0.445 0.997
Llama-13B LR 1.000 1.000 0.000
IMDB Zhou et al. (2021) 0.969 0.996 0.144
CE (Hendrycks & Gimpel, 2016) 0.961 0.995 0.206
TAPT (Gururangan et al., 2020) 0.965 0.995 0.159
SupCon (Khosla et al., 2020) 0.970 0.996 0.150
Uppaal et al. (2023) 0.990 0.998 0.012
Llama-7B LH 0.755 0.311 0.932
Llama-7B LR 1.000 1.000 0.001
Mistral-7B LH 0.767 0.943 0.926
Mistral-7B LR 0.999 0.998 0.003
Llama-13B LH 0.773 0.332 0.919
Llama-13B LR 1.000 1.000 0.000
Multi30K Zhou et al. (2021) 0.980 0.888 0.005
CE (Hendrycks & Gimpel, 2016) 0.962 0.920 0.175
TAPT (Gururangan et al., 2020) 0.956 0.922 0.167
SupCon (Khosla et al., 2020) 0.955 0.918 0.201
Uppaal et al. (2023) 1.000 1.000 0.000
Llama-7B LH 0.002 0.470 1.000
Llama-7B LR 1.000 1.000 0.000
Mistral-7B LH 0.002 0.470 1.000
Mistral-7B LR 0.995 0.998 0.008
Llama-13B LH 0.002 0.470 1.000
Llama-13B LR 1.000 1.000 0.000
NewsCategory Zhou et al. (2021) 0.955 0.969 0.383
CE (Hendrycks & Gimpel, 2016) 0.957 0.984 0.234
TAPT (Gururangan et al., 2020) 0.947 0.981 0.243
SupCon (Khosla et al., 2020) 0.962 0.986 0.219
Uppaal et al. (2023) 1.000 1.000 0.000
Llama-7B LH 0.014 0.128 1.000
Llama-7B LR 1.000 1.000 0.001
Mistral-7B LH 0.019 0.129 1.000
Mistral-7B LR 0.997 0.997 0.006
Llama-13B LH 0.017 0.128 0.999
Llama-13B LR 1.000 1.000 0.000
CLINC150 Zhou et al. (2021) 0.988 0.870 0.005
CE (Hendrycks & Gimpel, 2016) 0.964 0.844 0.189
TAPT (Gururangan et al., 2020) 0.959 0.830 0.213
SupCon (Khosla et al., 2020) 0.957 0.821 0.230
Uppaal et al. (2023) 1.000 1.000 0.000
Llama-7B LH 0.001 0.661 1.000
Llama-7B LR 1.000 1.000 0.000
Mistral-7B LH 0.001 0.661 1.000
Mistral-7B LR 0.995 1.000 0.008
Llama-13B LH 0.001 0.661 1.000
Llama-13B LR 1.000 1.000 0.000
Table 5: Results of far OOD detection, utilizing the same experimental setup as described by Uppaal et al. (2023). Results for methods not originating from our work are cited directly from Uppaal et al. (2023).
Ling No Llama-7B LH 0.552 0.215 0.934
Llama-7B LR 0.967 0.858 0.174
Llama-13B LH 0.534 0.182 0.929
Llama-13B LR 0.933 0.746 0.336
Yes NB 1.000 1.000 0.000
Logistic 1.000 1.000 0.000
KNN 0.968 0.932 0.021
SVM 1.000 1.000 0.000
XGBoost 0.995 0.973 0.017
LightGBM 0.997 0.979 0.008
RoBERTa 1.000 1.000 0.000
Spam-T5 1.000 1.000 0.000
Llama-7B LR 0.998 0.993 0.008
Llama-13B LR 0.997 0.988 0.012
SMS No Llama-7B LH 0.960 0.699 0.088
Llama-7B LR 0.866 0.582 0.487
Llama-13B LH 0.957 0.689 0.093
Llama-13B LR 0.810 0.518 0.761
Yes NB 0.988 0.949 0.113
Logistic 0.985 0.946 0.124
KNN 0.863 0.830 0.811
SVM 0.997 0.980 0.024
XGBoost 0.918 0.873 0.676
LightGBM 0.978 0.921 0.103
RoBERTa 0.997 0.988 0.004
Spam-T5 0.985 0.959 0.082
Llama-7B LR 1.000 1.000 0.000
Llama-13B LR 0.999 0.995 0.000
SpamAssassin No Llama-7B LH 0.964 0.884 0.096
Llama-7B LR 0.960 0.935 0.296
Llama-13B LH 0.956 0.897 0.169
Llama-13B LR 0.941 0.917 0.398
Yes NB 0.971 0.917 0.070
Logistic 0.992 0.986 0.029
KNN 0.931 0.935 0.578
SVM 0.990 0.983 0.046
XGBoost 0.994 0.989 0.019
LightGBM 1.000 0.999 0.000
RoBERTa 0.999 0.997 0.000
Spam-T5 0.996 0.994 0.012
Llama-7B LR 0.998 0.996 0.005
Llama-13B LR 0.994 0.989 0.019
Enron No Llama-7B LH 0.721 0.728 0.798
Llama-7B LR 0.991 0.989 0.043
Llama-13B LH 0.719 0.723 0.798
Llama-13B LR 0.992 0.990 0.035
Yes NB 0.992 0.991 0.035
Logistic 0.994 0.992 0.025
KNN 0.915 0.927 0.239
SVM 0.998 0.998 0.008
XGBoost 0.975 0.967 0.111
LightGBM 0.997 0.997 0.013
RoBERTa 1.000 1.000 0.001
Spam-T5 1.000 1.000 0.001
Llama-7B LR 0.999 0.999 0.001
Llama-13B LR 1.000 1.000 0.000
Table 6: Results of spam detection.
GSM8K SQUAD 2.0 7B $S_{q}$ 0.1116 0.0546 0.9894
$S_{a}$ 0.5463 0.0979 0.9947
$S_{q,a}$ 0.5363 0.0959 0.9924
$S_{a|q}$ 0.6877 0.1376 0.9704
13B $S_{q}$ 0.0519 0.0524 0.9955
$S_{a}$ 0.4958 0.0891 0.9955
$S_{q,a}$ 0.4017 0.0764 0.9947
$S_{a|q}$ 0.6144 0.1136 0.9765
BoolQ 7B $S_{q}$ 0.0538 0.1618 0.9955
$S_{a}$ 0.5045 0.2616 0.9932
$S_{q,a}$ 0.4797 0.2536 0.9879
$S_{a|q}$ 0.7156 0.4041 0.9310
13B $S_{q}$ 0.1008 0.1659 0.9924
$S_{a}$ 0.5507 0.2861 0.9833
$S_{q,a}$ 0.4778 0.2530 0.9795
$S_{a|q}$ 0.6967 0.3886 0.9303
PIQA 7B $S_{q}$ 0.9762 0.9779 0.0735
$S_{a}$ 0.9612 0.9746 0.0569
$S_{q,a}$ 0.9975 0.9983 0.0038
$S_{a|q}$ 0.9944 0.9944 0.0099
13B $S_{q}$ 0.9900 0.9906 0.0318
$S_{a}$ 0.8879 0.9331 0.1122
$S_{q,a}$ 1.0000 0.9999 0.0000
$S_{a|q}$ 1.0000 1.0000 0.0000
MATH SQUAD 2.0 7B $S_{q}$ 0.2139 0.2304 0.9474
$S_{a}$ 0.6384 0.3963 0.8916
$S_{q,a}$ 0.6527 0.4477 0.8436
$S_{a|q}$ 0.7385 0.5305 0.7914
13B $S_{q}$ 0.1880 0.2232 0.9460
$S_{a}$ 0.6098 0.3841 0.8890
$S_{q,a}$ 0.5628 0.3649 0.8882
$S_{a|q}$ 0.6786 0.4737 0.8166
BoolQ 7B $S_{q}$ 0.1303 0.4472 0.9658
$S_{a}$ 0.6135 0.7111 0.8580
$S_{q,a}$ 0.6361 0.7474 0.8008
$S_{a|q}$ 0.7507 0.8266 0.6870
13B $S_{q}$ 0.2612 0.5223 0.9304
$S_{a}$ 0.6488 0.7474 0.8148
$S_{q,a}$ 0.6384 0.7521 0.7818
$S_{a|q}$ 0.7350 0.8191 0.6854
PIQA 7B $S_{q}$ 0.9681 0.9902 0.0732
$S_{a}$ 0.9206 0.9775 0.1133
$S_{q,a}$ 0.9873 0.9962 0.0242
$S_{a|q}$ 0.9876 0.9956 0.0812
13B $S_{q}$ 0.9795 0.9938 0.0452
$S_{a}$ 0.8495 0.9572 0.1627
$S_{q,a}$ 0.9897 0.9969 0.0198
$S_{a|q}$ 0.9626 0.9895 0.0484
Table 7: Outcomes of OOD question detection in QA settings.

### v2

@@ 5 EXPERIMENTS
Yes Podolskiy et al., (2021) 0.998 0.994 0.008
SNIPS No Gangal et al., (2020) 0.955 0.903 0.192
Jin et al., (2022) 0.963 0.910 0.145
@@
Yes Podolskiy et al., (2021) 0.978 0.933 0.120
CLINC150 No Gangal et al., (2020) 0.883 0.677 0.463
Jin et al., (2022) 0.902 0.703 0.417
@@
Yes Podolskiy et al., (2021) 0.982 0.939 0.092
Table 2: Results for near OOD detection. Since the experimental configurations in the studies by Gangal et al., (2020), Podolskiy et al., (2021), and Jin et al., (2022) differ, we have replicated their methods and aligned the dataset splitting for consistency.
5.1 Far OOD Detection
For text data, detecting far OOD instances is relatively less challenging. Uppaal et al., (2023) have illustrated that utilizing the latent distance from a pretrained RoBERTa model significantly addresses far OOD detection, especially when the 20 Newsgroups (20NG) dataset serves as the in-distribution. In this context, we present that our proposed method also attains nearly perfect performance under the same experimental conditions, as detailed in Table 1.
Note that, the notation ‘Model-XB LH/LR’ is used, where ‘Model’ can be either Llama (Touvron et al.,, 2023) or Mistral (Jiang et al.,, 2023), denoting two popular open-source large language models (LLMs). Here, ‘Llama’ specifically refers to the Llama 2 version. The ‘XB’ indicates the model’s parameter count, either 7B or 13B. ‘LH’ stands for likelihood, signifying the use of $p_{\theta}(x)$ as the criterion for OOD detection; ‘LR’ denotes likelihood-ratio, referring to the employment of $S(x)$ as outlined in Equation (1) for OOD identification.
@@ 5.1 Far OOD Detection
Table 3: Results of spam detection. For the complete details, please see Appendix.
5.2 Near OOD Detection
We select the ROSTD (Gangal et al.,, 2020), SNIPS (Coucke et al.,, 2018), and CLINC150 (Larson et al.,, 2019) datasets for our near OOD detection experiments. The ROSTD and CLINC150 datasets are specifically crafted for OOD detection and include designated classes representing OOD data from the same domain. The SNIPS dataset comprises user utterances distributed among seven intent classes, such as GetWeather and RateBook. As it does not inherently provide OOD utterances, we classify the GetWeather and BookRestaurant intents as OOD for the purpose of our experiments. It’s noteworthy that this classification diverges from the one in the study by Jin et al., (2022), which does not explicitly detail their data splitting methodology.
Table 2 demonstrates that the likelihood ratio between the pretrained Llama model and the finetuned Llama model yields the highest performance among unsupervised OOD detection methods. In the case of CLINC150, the supervised OOD detection method introduced by Podolskiy et al., (2021) significantly surpasses our approach, a point that is further discussed in Section 6.
5.3 Spam Detection
Given that the concept of unsupervised OOD detection aligns closely with spam detection, we evaluate our method using the spam detection benchmark introduced by Labonne and Moran, (2023). This benchmark includes four specific spam detection datasets: Ling-Spam Dataset (Sakkis et al.,, 2003), SMS Spam Collection (Almeida et al.,, 2011), SpamAssassin Public Corpus, and Enron Email Dataset (Metsis et al.,, 2006). It compares the performance of both traditional and deep learning-based binary classifiers. Our method, being rooted in OOD detection, requires only the non-spam (ham) data for finetuning the LLM. Table 3 indicates that our method, without any spam data, can still reliably identify spam. Furthermore, when spam data is available, we can finetune the OOD proxy using this data and apply the likelihood ratio between the two finetuned LLMs. This approach achieves performance that is competitive with the SOTA in spam detection.
5.4 OOD Question Detection in QA Systems
We test the effectiveness of the criterion we introduced in Section 4 within a QA scenario. Here, we employ MetaMath (Yu et al.,, 2023), law-chat and medicine-chat (Cheng et al.,, 2023), which are Llama2 models finetuned on specific domains. The objective in this QA context is to identify OOD questions that originate from domains outside the model’s expertise. Filtering out OOD questions enhances the system’s robustness. For this evaluation, we designate the test sets of the GSM8K (Cobbe et al.,, 2021), MATH (Hendrycks et al.,, 2021), casehold (Zheng et al.,, 2021) and PubMedQA (Jin et al.,, 2019) datasets as In-D and use SQUAD (Rajpurkar et al.,, 2018), BoolQ (Clark et al.,, 2019), and PIQA (Bisk et al.,, 2020) as OOD datasets. Table 4 indicates that the criterion $S_{a}$ consistently outperforms the other evaluated criteria, with all its AUROC values exceeding 0.5. This demonstrates its effectiveness in detecting OOD. Notably, $S_{q}$ exhibits subpar performance in most scenarios, underscoring the necessity of our proposed approach that generates an answer and formulates the criterion based on the question-and-answer pair.
@@ 5.4 OOD Question Detection in QA Systems
MATH SQUAD 2.0 MetaMath-7B $S_{q}$ 0.2139 0.2304 0.9474
@@
casehold SQUAD 2.0 law-chat-7B $S_{q}$ 0.2463 0.2039 0.9987
$S_{a}$ 0.9082 0.8425 0.3717
$S_{q,a}$ 0.2048 0.1954 0.9989
$S_{a|q}$ 0.3915 0.2080 0.9992
BoolQ law-chat-7B $S_{q}$ 0.3869 0.5249 0.9985
$S_{a}$ 0.8878 0.9222 0.4753
$S_{q,a}$ 0.2692 0.4737 0.9988
$S_{a|q}$ 0.1307 0.3257 0.9996
PIQA law-chat-7B $S_{q}$ 0.9241 0.9717 0.3064
$S_{a}$ 0.9917 0.9978 0.0078
$S_{q,a}$ 0.8539 0.9440 0.4832
$S_{a|q}$ 0.0045 0.3743 0.9984
Table 4: OOD question detection in QA settings. For the complete details, please see Appendix.
@@ 6 DISCUSSIONS
Our far OOD detection experiments in Table 1 show that the AUROC for likelihood-based OOD detection methods can be exceptionally low. This finding echoes the phenomenon highlighted by Nalisnick et al., (2018), where language OOD data may unexpectedly exhibit higher likelihood values. Upon examining the characteristics of the datasets involved, it becomes apparent that the texts from the 20 Newsgroups (20NG) (Lang,, 1995) dataset are significantly longer than those from the comparative OOD datasets, such as SST-2 and RTE, especially in instances where the AUROC is notably low. This observation reveals a tendency among language models to assign higher likelihoods to shorter sentences, irrespective of their actual semantic content. It highlights the crucial importance of adopting the likelihood ratio rather than solely relying on raw likelihood values to enhance the accuracy of OOD detection.
@@
In the near OOD detection experiments presented in Table 2, Podolskiy et al., (2021)’s approach outperforms competing methods significantly. This superior performance may be attributed to the distinct class distribution across datasets. Specifically, the CLINC150 dataset comprises 150 In-D classes, in stark contrast to the SNIPS and ROSTD datasets, which offer a mere 7 (with only 5 for In-D) and 13 In-D classes, respectively. The substantially greater number of In-D classes in CLINC150 compared to the other datasets likely enhances Podolskiy et al., (2021)’s method’s ability to leverage the extensive class label information, thus yielding improved OOD detection results.
Some cases that LH is better than LR
In the spam detection experiments detailed in Table 3, particularly with data from the SMS and SpamAssassin datasets, we observe that without spam data, the likelihood (LH) method outperforms the likelihood ratio (LR). While LR generally shows superior and more consistent performance across most experiments — as LH can exhibit extremely poor performance in certain cases, a point elaborated in Section 6 — there are specific instances where LH is more effective.
7 RELATED WORKS
Building upon the OOD detection method using likelihood ratios introduced by Ren et al., (2019), Gangal et al., (2020) suggested employing the likelihood ratio between two LSTM language models. In their approach, one model functions as a “background model” representing OOD data and is trained on random combinations from the vocabulary.
Jin et al., (2022) used the likelihood ratio between a pretrained GPT-2 and a finetuned version of GPT-2 as a baseline to compare with their proposed contrastive learning-based OOD detection method. They showed that their proposed method can out perform the likelihood ratio based method. However, they only used GPT-2 in their likelihood ratio baseline, which is very small both in terms of training data and the model size by today standard.
In comparison, our study provides a more comprehensive analysis of likelihood ratio between base models and fine-tuned models using much larger LLMs. We show that leveraging these more advanced LLMs, likelihood ratio significantly outperforms results from Jin et al., (2022) in OOD detection. Additionally, in the current era, accessing and sharing both pretrained and finetuned LLMs has become considerably easier through online platforms such as Hugging Face. Using likelihood ratio in LLMs for OOD detection is easy, accessible, and very effective.
@@ 8 CONCLUDING REMARKS
ACKNOWLEDGEMENTS
In the early stages of this work, AZ was supported by a personal grant from Mrs. Yanshu Wu. In the middle and later stages, AZ was funded by the UKRI AI Hub in Generative Models (EP/Y028805/1). RB and TX were supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A. RB acknowledges funding by the German Research Foundation (DFG) for project 448588364 of the Emmy Noether Programme. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Tim Z. Xiao.